If one of the roots of the equation 2x^2 - x + C =0 is doubled the other root, find C.
Answer by ankor@dixie-net.com(22740)

If one of the roots of the equation 2x^2 - x + C = 0 is double the other root, find C.
We know that the roots have to be different signs, factors to
(2x+1)(x-1) FOILs to 2x^2 - 2x + x - 1 = 2x^2 - x - 1; c = -1
Roots
2x = -1
x = -.5
and
x = 1
: